# MOSA: Mixtures of Simple Adapters Outperform Monolithic Approaches in LLM-based Multilingual ASR

### v1

@@ III-D Ablation Study
A. Impact of the Number of Adapters. Table II presents the impact of varying the number of adapters on ASR performance. Regardless of whether 2, 3, 4, or 5 adapters are used, the average WER and monolingual WER consistently outperform the Ideal-LLM Base model. In other words, even with just two adapters (60% training parameters of Ideal-LLM Base), the performance surpasses that of Ideal-LLM Base. As the number of adapters increases, the average WER gradually decreases, and the performance across most individual languages also improves. However, using five adapters does not yield further improvement, likely due to the model becoming too large relative to the available training data. Furthermore, we investigated the performance of using a single adapter without incorporating a router. The results indicate that, across multiple languages, the single-adapter setup consistently underperforms compared to the multi-adapter configuration. This is particularly evident in low-resource languages such as Polish. The relatively strong performance on Portuguese may be attributed to the adapter being optimized for that specific language at checkpointing. To further examine this, we fine-tuned the single-adapter model individually on the data of each language, resulting in eight distinct systems (denoted as $1^{*}$). However, the improvements observed were still marginal across all languages. These findings suggest that multiple adapters are capable of capturing language-specific or shared knowledge, thereby alleviating the burden on a single adapter and leading to better overall results.
Figure 2: T-SNE visualization of aligned speech embeddings.
B. Analysis of Speech Embeddings Across Languages. We conducted a distributional analysis of the speech embeddings generated by the projector of MOSA-Base. Specifically, we applied t-SNE[36] visualization on the test set of the MLS dataset, as shown in Figure 2. The results reveal that the embeddings from all eight languages in MLS form clearly separable clusters, with only a few outlier utterances. This indicates that the model is capable of treating each language similarly to a monolingual system. These findings further demonstrate that mixing multiple adapters can better accommodate multilingual settings.
Figure 3: Adapter weight distribution across languages.
@@
IV Conclusion
In this work, we observe that a single, complex adapter struggles to effectively adapt across multiple languages. To address this, we propose MOSA (Mixture of Simple Adapters), a method based on the MoE mechanism that combines multiple lightweight adapters. This design allows different experts to learn either language-specific or shared knowledge, enabling high-resource languages to benefit low-resource ones and alleviating the data scarcity problem in multilingual settings. Moreover, this approach simplifies the overall model architecture. Experimental results indicate that MOSA-Base achieves a 15.4% relative improvement in average WER over Baseline-Base and consistently outperforms it across all evaluated languages. Notably, MOSA-Base maintains superior performance even when trained with only 60% of the parameters used by Baseline-Base. Likewise, MOSA-Large surpasses Baseline-Large in average WER and exhibits enhanced robustness to data imbalance. Ablation studies further reveal that MOSA is more effective at handling individual languages and at learning both language-specific and shared linguistic features. These results suggest that, in LLM-based ASR, using a mixture of simple adapters is more effective than relying on a single, complex adapter architecture.

### v2

@@ 3.4 Ablation Study
A. Impact of the Number of Adapters. Table 2 presents the impact of varying the number of adapters on ASR performance. Regardless of whether 2, 3, 4, or 5 adapters are used, the average WER and monolingual WER consistently outperform the Ideal-LLM Base model. In other words, even with just two adapters (60% training parameters of Ideal-LLM Base), the performance surpasses Ideal-LLM Base by 13.3%. As the number of adapters increases, the average WER gradually decreases, and the performance across most individual languages also improves. However, using five adapters does not yield further improvement, likely due to the model becoming too large relative to the available training data. Furthermore, we investigated the performance of using a single adapter without incorporating a router. The results indicate that, across multiple languages, the single-adapter setup consistently underperforms compared to the multi-adapter configuration. This is particularly evident in low-resource languages such as Polish. The relatively strong performance on Portuguese may be attributed to the adapter being optimized for that specific language at checkpointing. To further examine this, we fine-tuned the single-adapter model individually on the data of each language, resulting in eight distinct systems (denoted as $1^{*}$). However, the improvements observed were still marginal across all languages. These findings suggest that multiple adapters are capable of capturing language-specific or shared knowledge, thereby facilitating positive transfer from high-resource to low-resource languages.
Fig. 2: T-SNE visualization of aligned speech embeddings.
B. Analysis of Speech Embeddings Across Languages. We conducted a distributional analysis of the speech embeddings generated by the projector of MOSA-Base. Specifically, we applied t-SNE [29] visualization on the test set of the MLS dataset, as shown in Figure 2. The results reveal that the embeddings from all eight languages in MLS form clearly separable clusters, with only a few outlier utterances. This indicates that the model is capable of treating each language similarly to a monolingual system, thereby mitigating multilingual parameter interference. These findings further demonstrate that mixing multiple adapters can better accommodate multilingual settings.
Fig. 3: Adapter weight distribution across languages.
@@
4 Conclusion
In this work, we observe that a single, complex projector struggles to effectively align representations across multiple languages. To address this, we propose MOSA within the projector, which combines multiple lightweight adapters. This design enables different experts to capture either language-specific or shared knowledge, thereby mitigating multilingual parameter interference and facilitating positive transfer from high-resource to low-resource languages. Experimental results indicate that MOSA-Base achieves a 15.4% relative improvement in average WER over Ideal-LLM Base, consistently outperforming it across all evaluated languages. Notably, MOSA maintains superior performance even when trained with only 60% of the parameters used by Ideal-LLM Base, demonstrating its exceptional parameter efficiency and robustness to data imbalance.
5 Acknowledgment
This work was supported by the Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100).